OpenStudy (anonymous):

Give the formula for the arithmetic sequence. 1, -1, -3, -5, ... Answer tn = 1 + 2(n – 1) tn = 1 – 2(n – 1) tn = 1 + 2(n – 1) tn = 1 – 2(n + 1)

OpenStudy (anonymous):

here, first term =1 common difference =t2-t1=-1-1=-2

OpenStudy (anonymous):

AND tn = a+(n-1)d

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

third option